Station 1 Station 2 Station 3 Station 4 Station 5 Station 6 Arrivals PT = 1.5 hour 0.2 Parts/hr T WIP = 25pcs 2 hour PT 3.5 hour PT = 2.5 hour PT 1.25 hour PT 1.25 hour

Which Station is the Bottleneck?

What is the Bottleneck Rate (BNR)?

What is the Raw Processing Time (RPT)?

What is the Critical WIP level?

Best-Case Throughput?

Best-Case Cycle Time?

Worst-Case Throughput?

Worst-Case Cycle Time?

Practical Worst-Case Throughput?

Practical Worst-Case Cycle Time?   

If Actual Throughput is 0.18, in what region are we operating?

What suggestion would you make to improve this line flow, explain?
Answer #1

Answer 1: station 2 is bottle neck, because it takes 3.5 hours for the operation

Answer 2: BNR= 1/3.5= 0.28 units per hour

Answer 3: Raw processing time (RPT)= total time to complete 1 unit= 12 hours

Answer 4: critical WIP= BNR*RPT= 3.42

Answer 5: best throughput= critical WIP/RPT= BNR i.e 0.28 units per hour

Answer 6: best case cycle time= RPT = 12 hours

Note: Chegg's guideline mandate to solve first four subparts of the question